# Quantification of Interdependent Emotion Dynamics in Online Interactions

### v2

@@ Abstract
Figure 1: (a) Screenshot of a YouTube live video on the Theranos scandal involving Elizabeth Holmes with a live chat section on the right. The video content is captured by transcripts (subtitles), which we use to proxy exogenous emotional stimuli for the live chats. The live chat section displays timestamped messages from users reacting to the video content in real-time as the live video streams. We highlight an example of a transcript labeled as sad in blue, and a live chat message labeled as angry in red. (b,c) We visualize the extraction of emotions from the video and live chat in the screenshot above. We plot the arrival of live chat messages that are labeled as sad (angry), indicated with diamond markers in blue (red). The shared x-axis shows the time in the video in units of minutes. The y-axis is unit-less. We label emotions non-exclusively. For instance, the sentence “Holmes is a victim of the fake news media.” is labeled as both sad and angry. We assume these emotions are generated by the latent, inhomogeneous intensity defined by expression (1). (d) Time-varying component of the emotion sad in the video (transcript). Blue dots annotate the arrival of transcripts that are labeled as sad. The black dashed line represents the temporal function $S^{\text{sad}}(t)$, capturing the presence of sad emotions within the video.
@@
Using this framework, we analyze the joint dynamics of six basic emotions in YouTube live chats. We classify emotions originating from video content as exogenous (exo) inputs and those expressed in the chat as endogenous (endo) responses driven by peer interaction. By estimating the parameters of a multivariate Hawkes model across 397 videos totaling more than 780 hours of content, we show that user emotional expressions are influenced by endogenous interactions roughly four times more than by exogenous video content. This suggests that in online environments, emotions are not simply reactions to external stimuli but are actively sustained and amplified through social feedback.
@@ Methods
We examine the dynamics of emotions in online discussions through a collection of YouTube Live videos. YouTube is a video sharing platform that caters to a wide range of users across different cultures. YouTube Live allows the audience to chat with each other while the video is being played in real time. We refer to live discussions as live chats and a single user live comment as a live chat message, or just a message. Videos with the live chat replay function record time-stamped live discussions and allow viewers to retrieve the live experience by simultaneously replaying the video content and live chat messages. Figure 1(a) shows an example of a YouTube live video (left) with a live chat section (right) where viewers post messages visible to the entire audience as the video streams.
We examine emotion dynamics in YouTube live videos with replay functionality across 27 topics ranging from comedy and documentary to sports and travel (see SI Appendix for the complete list). We first compile the list of topic keywords and collect completed live video IDs for each topic keyword using the keyword search function via the YouTube API in Python. 111 Code and data are available at https://github.com/ivylyis/Quantification-of-the-Self-Excited-Emotion-Dynamics-in-Online-Interactions. We initially extract both the video transcripts and live chat messages associated with each video. This process yields two distinct lists: one of time-stamped chat messages and another of time-stamped video transcripts (subtitles). We proceed by selecting only those videos in which the median interval between successive live chat messages ranges from 1 second to 5 minutes. Additionally, we disregard videos where more than 70% of the chat messages are in a language other than English. For the videos retained, we eliminate non-English live chat messages. Furthermore, we delete any chat messages that precede the first or succeed the last transcript time in each video. Ultimately, this results in a collection of 673,551 live chat messages from 1,957 videos.
@@
Multiple emotions can arise simultaneously in a single sentence [39, 41]. We thus assign to each text a vector of length 6 indicating the non-exclusive presence of each of the six emotions. This amounts to a multilabel classification problem. For training data, we rely on the SemEval-2018 dataset containing 6,838 Tweets with non-exclusive emotion labels across 11 emotions [53]. We group semantically related emotions into six core categories by mapping anticipation, optimism, love, and trust to joy; pessimism to sadness; and retaining sadness, anger, disgust, fear, and surprise in their original categories. Subsequently, we fine-tune a Roberta transformer model [54]. The transformer takes as input a text and gives as output the probabilities for each of the 6 basic emotions via 6 independent sigmoid activations. We use the binary cross-entropy loss function for training. During prediction, we convert each sigmoid output to a binary indicator at the 0.5 cutoff. This allows us to assign to each live chat message and video subscript a subset of the emotions $\mathcal{E}$ (see Figure 1 for an example).
For every video and emotion analyzed, we determine the average count of live chat messages per minute. To exclude outliers or inconsistencies in live chat data, only those videos where this average falls between the $20^{\text{th}}$- and $80^{\text{th}}$-quantiles of all videos are included. This process results in a compiled set of 92,412 live chat messages that span 397 videos, totaling 780 hours. The median video length is 107 minutes and per video, the median number of live chat messages is 204. When considering all videos, the median number of live chat messages per minute is as follows: 1.63 for joy, 0.17 for disgust, 0.16 for anger, 0.11 for sadness, and for both fear and surprise, 0.03. Additional details are provided in the SI Appendix.
@@
Emotional contagion, the phenomenon in which emotions spread from one individual to another, is naturally captured by the self-exciting nature of the Hawkes process [55]. In our context, emotions are modeled as events whose statistical occurrence follows a point process. Because we have more than just one emotion and the emotions (presumably) cross-excite each other, we represent the dynamics of $|\mathcal{E}|=6$ distinct emotions using a multivariate Hawkes process [65]. It is our goal to understand to what extent a given emotion $f\in\mathcal{E}$ triggers another emotion $e\in\mathcal{E}$ either exogenously from the video to the live chat or endogenously from live chats to live chats. Here, the adjectives exogenous and endogenous refer to the point of view of live chat: If emotion $f$ appears in the video and triggers an emotion $e$ in the live chat, we call it exogenous (exo). In contrast, if emotion $f$ also appears in the live chat and subsequently triggers emotion $e$ in the live chat, we call it endogenous (endo). We assume that emotions in the video and in the live chat can trigger emotions in the live chat, but that emotions in the live chat cannot influence emotions in the video. This is precisely the case if the video is pre-recorded (e.g., a documentary, as in Figure 1), and approximately the case where the video involves a live streamer who could potentially interact with the viewers.
@@
We model the arrival of emotions as a multivariate Hawkes process given by expression (1). For emotion $e\in\mathcal{E}$, within a given live chat session, we observe $N^{e}$ events, with the $j^{\text{th}}$ event taking place at time $t_{j}^{e}$. We define $t_{0}^{e}\equiv 0$, $t_{N^{e}+1}^{e}\equiv T$ where $T$ is the duration of the video. In this way, the observation period for the entire process is $\{t_{i}^{e}\mid i=0,\ldots,N^{e}+1\}$. Similarly, we denote by $\tau^{e}_{i}$ the time at which the $i$-th subscript of emotion $e$ appears in the video. We measure time in units of minutes. In the rest of this subsection, we describe the process of parameterizing the time-varying exogenous video influence $S^{f}(t)$. In essence, $S^{f}(t)$ denotes the level of emotion $f$ exhibited in the video at a particular time $t$ (refer to Figure 1(d) for an illustration). Readers who are mainly interested in the qualitative results of our study may skip this and the next subsection and continue with the Results Section.
Previous studies have shown that assuming the exogenous influence $\mu^{e}(t)$ to be constant can lead to false attribution to endogenous effects [66, 60] (see SI Appendix for additional confirmation). The YouTube live videos provide us with the unique opportunity to observe a time-varying influence of video content on user discussions. We use the video transcript to capture the time-varying emotional content of the video. For example, it is reasonable to assume that a particularly sad scene in the video induces sad emotions in the live chat. In addition, cross-influence can be expected; for example, a sad scene in the video induces angry emotions in the live chat (Figure 1). In particular, this is to be distinguished from a sad scene in the video triggering a sad message in the live chat, which in turn triggers either a sad or angry message. To capture such cross-influence from the videos, we parameterize the exogenous intensity of emotion $e$ as
@@
To fix $\mu$ and $\sigma$, we assume that the intensity of each transcript peaks $2$ seconds after appearance and that 50% of the emotion intensity for each transcript is manifested within $10$ seconds of transcript appearance. In other words, we assume that the maximum of the log-normal function lies at $2$ seconds, $\exp(\mu-\sigma^{2})=2$, and that the median is equal to $10$ seconds, $\exp(\mu)=10$. Numerically solving for $\mu$ and $\sigma$ yields $2.3$ and $1.3$ respectively in units of minutes. To ensure that our results are not strongly dependent on this choice of $\mu$ and $\sigma$, we have checked that our results remain qualitatively similar for different values and functional shapes (SI Appendix). Alternatively, these parameters could be directly fitted from the data, a task which we leave for future research.
@@
where $M^{f}_{1}\equiv\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {S^{f}(s)}$ is a constant that we calculate numerically. A detailed derivation of Equation (5) is found in the SI Appendix.
@@
Note that Equation (5) represents the log-likelihood for a single video, $\log L^{e}=\log L^{e}_{k}$ where $k$ enumerates the videos in our data sample. Thus, we aggregate $\log L^{e}_{k}$ across all videos and estimate the model parameters by maximizing $\sum_{k}\log L^{e}_{k}$. The total number of parameters is $142|\mathcal{E}|+2=$ 14 per emotion. This amounts to a total of $|\mathcal{E}|\left(2|\mathcal{E}|+2\right)=84$ parameters to estimate for $|\mathcal{E}|=6$ emotions. However, the parameters for each emotion can be estimated independently because there is no interdependence between the parameters between different types of emotions. This gives rise to six independent fits of $14$ parameters each. To verify that the obtained fits are valid, we first test our model with synthetically generated data and obtain reliable performance (see SI Appendix). We also refer to the SI Appendix for additional robustness checks with respect to different parameterizations and subsets of emotions. To mitigate potential biases, we estimate our model parameters using bootstrapped samples. We bootstrap our data $10$ times, each time sampling only 60% of all available videos. This allows us to estimate the parameters as averages across bootstrapped samples, and we use standard deviations as error bars. The fitted parameters are shown in Figure 2. These parameters are interpreted in the Results Section and discussed more qualitatively in the Discussion Section.
@@ Results
Figure 2: Visualizations of estimated parameters $\nu^{e,f}$, $\mu^{e}_{0}$, $\alpha^{e,f}$, and $\gamma^{e}$ of the intensity defined by expression (1) from the maximization of the log-likelihood function (5). The parameters are fitted ten times, each time sampling 60% (238) of the total 397 videos at random. Values and error bars are then obtained as sample mean and standard deviations across all 10 fits, respectively. Entry $\alpha^{e,f}$ of the $\alpha$ matrix represents excitation from emotion $f$ to emotion $e$ in the live chat, while $\gamma^{e}$ represents the characteristic time-scale over which past emotions trigger new emotions $e$. The $\nu$ matrix demonstrates how emotions in the video trigger emotions in the live chat, while $\mu_{0}$ illustrates the spontaneous baseline intensity of spontaneous emotion expression.
@@
Videos provide a shared experience among viewers, resulting in increased emotional intensities during live broadcast events [12]. However, it is unclear to what extent the emotional experience and amplification stem from the live videos themselves or the interactions among peers [12]. The Hawkes framework (1) allows us to answer this question by comparing exogenously and endogenously triggered events. The exogenous component, $\mu^{e}(t)=\mu^{e}_{0}+\mu_{1}^{e}(t)$, represents influences from video content and elsewhere. The endogenous component, $\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}<t}\phi^{e,f}\left(t-t_{j}^{f}\right)$, captures the influence of previous live chat messages and is therefore derived from interactions with peers. At any given time, we can determine the fraction of exo-influence as $R^{e}_{\text{exo}}(t)=\left.\mu^{e}(t)\right/\lambda^{e}(t)$ and the ratio of endo influence as $R^{e}_{\text{endo}}(t)=1-R^{e}_{\text{exo}}(t)$ where all quantities are calculated with the estimated parameters (shown in Figure 2 and further discussed below). Subsequently, we calculate the average exo- and endo-ratios, $\langle R^{e}_{\text{exo}}\rangle$ and $\langle R^{e}_{\text{endo}}\rangle$, of a given video by averaging over time. To put it differently, for each video, $\langle R^{e}_{\text{exo (exo)}}\rangle$ represents the mean portion of intensity for emotion $e$ that is elicited exogenously (endogenously). Figure 3 (left) shows the distribution of these quantities over all videos in our sample. We observe that despite the constant exogenous influence from video feeds, endogenous influence dominates the emotion dynamics across emotion types. This shows that emotional expressions in live discussions are disproportionately driven by social interactions with active participants rather than passive consumption of video content. In particular, we see that the emotion of joy has the highest ratio of endogenous influence, being 80% driven by prior participants. In other words, joy is four times more likely to be triggered endogenously than exogenously. In contrast, and according to common sense, the emotion of surprise is the most responsive to exogenous stimuli, where roughly 50% of the total intensity is triggered exogenously.
@@
Similarly, we can calculate the ratio of purely exogenous spontaneous emotions relative to video-induced emotions, that is $R^{e}_{0}=\mu^{e}_{0}/\left(\mu^{e}_{0}+\mu^{e}_{1}(t)\right)$ and its complement $R^{e}_{1}=1-R^{e}_{0}.$ Figure 3 (right) shows the distribution of $\langle R^{e}_{0}\rangle$ versus $\langle R^{e}_{1}\rangle$ over the set of all videos. The rate of chats induced by videos is typically one-fourth the rate of spontaneous chats. We note, however, that there is significant heterogeneity across video categories (SI Appendix). In summary, this highlights that live chat discussions elicit emotional expressions largely due to social dynamics rather than simply watching video content. In the SI Appendix, we further support these findings with additional robustness checks, specifically focusing on outcomes when fitting only selected subsets of emotions or video categories.
@@
Across our 10 bootstrapped estimates, we estimate the spectral radius to be $0.80\pm 0.008$, which places the system well below criticality. This results in an expected cascade size of $5.0\pm 0.2$, meaning that each exogenous emotion, on average, generates a cascade of five total events before activity dies out. Thus, while the system exhibits substantial self-excitation, it remains far from the near-critical regime. A similar pattern holds at the level of individual emotions, where the rate of endogeneity for each emotion type $f$ is given by $\sum_{e}\alpha^{e,f}$.

### v3

@@ Introduction
Figure 1: (a) Screenshot of a YouTube live video on the Theranos scandal involving Elizabeth Holmes with a live chat section on the right. The video content is captured by transcripts (subtitles), which we use to proxy exogenous emotional stimuli for the live chats. The live chat section displays timestamped messages from users reacting to the video content in real-time as the live video streams. We highlight an example of a transcript labeled as sad in blue, and a live chat message labeled as angry in red. (b,c) We visualize the extraction of emotions from the live chat in the screenshot above. We plot a subset of live chat messages from the video sample that are labeled as sad (angry), indicated with diamond markers in blue (red). The shared x-axis shows the time in the video in units of minutes. The y-axis is unit-less. We label emotions non-exclusively. For instance, the sentence “Holmes is a victim of the fake news media.” is labeled as both sad and angry. We assume these emotions are generated by the latent, inhomogeneous intensity defined by expression (1). (d) Time-varying component of the emotion sad in the video (transcript). Blue dots annotate the arrival of transcripts that are labeled as sad, shown with a 2-second rightward shift for alignment with peaks in the signal for visual convenience (see Methods section for details). The black dashed line represents the temporal function $S^{\text{sad}}(t)$, capturing the presence of sad emotions within the video.
@@
Using this framework, we analyze the joint dynamics of six basic emotions in YouTube live chats. We define the system of interest as the collective set of viewers participating in the live chat. Under this group-level framework, emotional expressions triggered by prior messages within the chat are considered endogenous, while influences originating outside the chat, primarily from the video content, are treated as exogenous. This distinction allows us to study how social interaction shapes group-level emotion dynamics during livestreams. Concretely, we classify emotions originating from video content as exogenous (exo) inputs and those expressed in the chat as endogenous (endo) responses driven by peer interaction. By estimating the parameters of a multivariate Hawkes model across 397 videos totaling more than 780 hours of content, we show that user emotional expressions are influenced by endogenous interactions roughly four times more than by exogenous video content. This suggests that in online environments, emotions are not simply reactions to external stimuli but are actively sustained and amplified through social feedback.
@@ Methods
We examine the dynamics of emotions in online discussions through a collection of YouTube Live videos. YouTube is a video-sharing platform that caters to a wide range of users across different cultures. YouTube Live allows the audience to chat with each other while the video is being played in real time. We refer to live discussions as live chats and a single user live comment as a live chat message, or just a message. Videos with the live chat replay function record time-stamped live discussions and allow viewers to retrieve the live experience by simultaneously replaying the video content and live chat messages. Figure 1(a) shows an example of a YouTube live video (left) with a live chat section (right) where viewers post messages visible to the entire audience as the video streams.
We examine emotion dynamics in YouTube live videos with replay functionality across 27 topics ranging from comedy and documentary to sports and travel (see Appendix A for the complete list). We first compile the list of topic keywords and collect completed live video IDs for each topic keyword using the keyword search function via the YouTube API in Python. 111 Code and data are available at https://github.com/ivylyis/Quantification-of-the-Self-Excited-Emotion-Dynamics-in-Online-Interactions. We initially extract both the video transcripts and live chat messages associated with each video. This process yields two distinct lists: one of time-stamped chat messages and another of time-stamped video transcripts (subtitles). Notably, YouTube Live Chat offers two display modes: an unfiltered chronological stream (“Live Chat”) and a filtered version (“Top Chat”) that removes spam and inappropriate messages 222 https://support.google.com/youtube/answer/15268877?hl=en-GB . We use the “Live Chat” version in our data collection process, ensuring that all messages are collected without algorithmic filtering. While viewers can choose different display modes when watching the video, we note that YouTube’s filtering is not personalized and applies uniformly across all viewers. This stands in contrast to the personalized recommendation algorithms commonly used on other social media platforms, which can introduce more substantial distortions in the content users are exposed to.
We proceed by selecting only those videos in which the median interval between successive live chat messages ranges from 1 second to 5 minutes. Additionally, we disregard videos where more than 70% of the chat messages are in a language other than English. For the videos retained, we eliminate non-English live chat messages. Furthermore, we delete any chat messages that precede the first or succeed the last transcript time in each video. Ultimately, this results in a collection of 673,551 live chat messages from 1,957 videos.
@@
Multiple emotions can arise simultaneously in a single sentence [39, 41]. We thus assign to each text a vector of length 6 indicating the non-exclusive presence of each of the six emotions. This amounts to a multilabel classification problem. For training data, we rely on the SemEval-2018 dataset containing 6,838 Tweets with non-exclusive emotion labels across 11 emotions [53]. We group semantically related emotions into six core categories by mapping anticipation, optimism, love, and trust to joy; pessimism to sadness; and retaining sadness, anger, disgust, fear, and surprise in their original categories. Subsequently, we fine-tune a Roberta transformer model [54]. The transformer takes as input a text and gives as output the probabilities for each of the 6 basic emotions via 6 independent sigmoid activations. We retain emojis in the text input, as they serve as an important channel for emotional expression in online communication. We use the binary cross-entropy loss function for training. During prediction, we convert each sigmoid output to a binary indicator at the 0.5 cutoff. This allows us to assign to each live chat message and video subscript a subset of the emotions $\mathcal{E}$. We provide examples of live chat messages and the corresponding assigned emotion labels in Appendix A.
For every video and emotion analyzed, we determine the average count of live chat messages per minute. To exclude outliers or inconsistencies in live chat data, only those videos where this average falls between the $20^{\text{th}}$- and $80^{\text{th}}$-quantiles of all videos are included. This process results in a compiled set of 92,412 live chat messages that span 397 videos, totaling 780 hours. The median video length is 107 minutes and per video, the median number of live chat messages is 204. When considering all videos, the median number of live chat messages per minute is as follows: 1.63 for joy, 0.17 for disgust, 0.16 for anger, 0.11 for sadness, and for both fear and surprise, 0.03. Additional details are provided in Appendix A.
@@
Emotional contagion, the phenomenon in which emotions spread from one individual to another, is naturally captured by the self-exciting nature of the Hawkes process [55]. In our context, emotions are modeled as events whose statistical occurrence follows a point process. Because we have more than just one emotion and the emotions (presumably) cross-excite each other, we represent the dynamics of $|\mathcal{E}|=6$ distinct emotions using a multivariate Hawkes process [65]. It is our goal to understand to what extent a given emotion $f\in\mathcal{E}$ triggers another emotion $e\in\mathcal{E}$ either exogenously from the video to the live chat or endogenously from live chats to live chats. Here, the adjectives exogenous and endogenous refer to the point of view of live chat: If emotion $f$ appears in the video and triggers an emotion $e$ in the live chat, we call it exogenous (exo). In contrast, if emotion $f$ also appears in the live chat and subsequently triggers emotion $e$ in the live chat, we call it endogenous (endo). We assume that emotions in the video and in the live chat can trigger emotions in the live chat, but that emotions in the live chat cannot influence emotions in the video. We demonstrate in Appendix E that our results remain qualitatively consistent both with and without the subset of videos for which this assumption does not necessarily hold.
@@
We model the arrival of emotions as a multivariate Hawkes process given by expression (1). For emotion $e\in\mathcal{E}$, within a given live chat session, we observe $N^{e}$ events, with the $j^{\text{th}}$ event taking place at time $t_{j}^{e}$. We define $t_{0}^{e}\equiv 0$, $t_{N^{e}+1}^{e}\equiv T$ where $T$ is the duration of the video. In this way, the observation period for the entire process is $\{t_{i}^{e}\mid i=0,\ldots,N^{e}+1\}$. Similarly, we denote by $\tau^{e}_{i}$ the time at which the $i$-th subscript of emotion $e$ appears in the video. We express time in units of minutes throughout the analysis and event times are treated as continuous variables throughout the model fitting process. The live chat messages have a temporal resolution measured in milliseconds while the video transcripts are measured in centiseconds. For example, an event occurring at 5 minutes and 3.2 seconds is recorded at 5.0533.
In the remainder of this subsection, we describe the process of parameterizing the time-varying exogenous video influence $S^{f}(t)$. In essence, $S^{f}(t)$ denotes the level of emotion $f$ exhibited in the video at a particular time $t$ (refer to Figure 1(d) for an illustration). Readers who are mainly interested in the qualitative results of our study may skip this and the next subsection and continue with the Results Section.
Previous studies have shown that assuming the exogenous influence $\mu^{e}(t)$ to be constant can lead to false attribution to endogenous effects [66, 60] (see Appendix B for additional confirmation). The YouTube live videos provide us with the unique opportunity to observe a time-varying influence of video content on user discussions. We use the video transcript to capture the time-varying emotional content of the video. For example, it is reasonable to assume that a particularly sad scene in the video induces sad emotions in the live chat. In addition, cross-influence can be expected; for example, a sad scene in the video induces angry emotions in the live chat (Figure 1). In particular, this is to be distinguished from a sad scene in the video triggering a sad message in the live chat, which in turn triggers either a sad or angry message. To capture such cross-influence from the videos, we parameterize the exogenous intensity of emotion $e$ as
@@
To fix $\mu$ and $\sigma$, we assume that the intensity of each transcript peaks $2$ seconds after appearance and that 50% of the emotion intensity for each transcript is manifested within $10$ seconds of transcript appearance. In other words, we assume that the maximum of the log-normal function lies at $2$ seconds, $\exp(\mu-\sigma^{2})=2$, and that the median is equal to $10$ seconds, $\exp(\mu)=10$. Numerically solving for $\mu$ and $\sigma$ yields $2.3$ and $1.3$ respectively in units of minutes. To ensure that our results are not strongly dependent on this choice of $\mu$ and $\sigma$, we have checked that our results remain qualitatively similar for different values and functional shapes (see Appendix E). Alternatively, these parameters could be directly fitted from the data, a task which we leave for future research.
@@
where $M^{f}_{1}\equiv\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {S^{f}(s)}$ is a constant that we calculate numerically. A detailed derivation of Equation (5) is found in Appendix C.
@@
Note that Equation (5) represents the log-likelihood for a single video, $\log L^{e}=\log L^{e}_{k}$ where $k$ enumerates the videos in our data sample. Thus, we aggregate $\log L^{e}_{k}$ across all videos and estimate the model parameters by maximizing $\sum_{k}\log L^{e}_{k}$. The total number of parameters is $142|\mathcal{E}|+2=$ 14 per emotion. This amounts to a total of $|\mathcal{E}|\left(2|\mathcal{E}|+2\right)=84$ parameters to estimate for $|\mathcal{E}|=6$ emotions. However, the parameters for each emotion can be estimated independently because there is no interdependence between the parameters between different types of emotions. This gives rise to six independent fits of $14$ parameters each. To verify that the obtained fits are valid, we first test our model with synthetically generated data and obtain reliable performance (see Appendix D). We also refer to Appendix E for additional robustness checks with respect to different parameterizations and subsets of emotions. To mitigate potential biases, we estimate our model parameters using bootstrapped samples. We bootstrap our data $100$ times, each time sampling only 60% of all available videos. This allows us to estimate the parameters as averages across bootstrapped samples, and we use standard deviations as error bars. The fitted parameters are shown in Figure 2. These parameters are interpreted in the Results Section and discussed more qualitatively in the Discussion Section.
@@ Results
Figure 2: Visualizations of estimated parameters $\nu^{e,f}$, $\mu^{e}_{0}$, $\alpha^{e,f}$, and $\gamma^{e}$ of the intensity defined by expression (1) from the maximization of the log-likelihood function (5). The parameters are fitted ten times, each time sampling 60% (238) of the total 397 videos at random. Values and error bars are then obtained as sample mean and standard deviations across all 100 fits, respectively. Entry $\alpha^{e,f}$ of the $\alpha$ matrix represents excitation from emotion $f$ to emotion $e$ in the live chat, while $\gamma^{e}$ represents the characteristic time-scale over which past emotions trigger new emotions $e$. The $\nu$ matrix demonstrates how emotions in the video trigger emotions in the live chat, while $\mu_{0}$ illustrates the spontaneous baseline intensity of spontaneous emotion expression. Note that the $\gamma^{e}$ and $\mu_{0}^{e}$ columns share the same $y$-axis emotion labels as the $\alpha$ and $\nu$ matrices, respectively.
@@
Videos provide a shared experience among viewers, resulting in increased emotional intensities during live broadcast events [12]. However, it is unclear to what extent the emotional experience and amplification stem from the live videos themselves or the interactions among peers [12]. The Hawkes framework (1) allows us to answer this question by comparing exogenously and endogenously triggered events. The exogenous component, $\mu^{e}(t)=\mu^{e}_{0}+\mu_{1}^{e}(t)$, represents influences from video content and elsewhere. The endogenous component, $\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}<t}\phi^{e,f}\left(t-t_{j}^{f}\right)$, captures the influence of previous live chat messages and is therefore derived from interactions with peers. At any given time, we can determine the fraction of exo-influence as $R^{e}_{\text{exo}}(t)=\left.\mu^{e}(t)\right/\lambda^{e}(t)$ and the ratio of endo influence as $R^{e}_{\text{endo}}(t)=1-R^{e}_{\text{exo}}(t)$ where all quantities are calculated with the estimated parameters (shown in Figure 2 and further discussed below). Subsequently, we calculate the average exo- and endo-ratios, $\langle R^{e}_{\text{exo}}\rangle$ and $\langle R^{e}_{\text{endo}}\rangle$, of a given video by averaging over time. To put it differently, for each video, $\langle R^{e}_{\text{exo (exo)}}\rangle$ represents the mean portion of intensity for emotion $e$ that is elicited exogenously (endogenously). Figure 3 (left) shows the distribution of these quantities over all videos in our sample. We observe that despite the constant exogenous influence from video feeds, endogenous influence dominates the emotion dynamics across emotion types. This shows that emotional expressions in live discussions are disproportionately driven by social interactions with active participants rather than passive consumption of video content. In particular, we see that the emotion of joy has the highest ratio of endogenous influence, being on average 75% driven by prior participants. In other words, joy is three times more likely to be triggered endogenously than exogenously. In contrast, and according to common sense, the emotion of surprise is the most responsive to exogenous stimuli, where roughly 50% of the total intensity is triggered exogenously.
@@
Similarly, we can calculate the ratio of purely exogenous spontaneous emotions relative to video-induced emotions, that is $R^{e}_{0}=\mu^{e}_{0}/\left(\mu^{e}_{0}+\mu^{e}_{1}(t)\right)$ and its complement $R^{e}_{1}=1-R^{e}_{0}.$ Figure 3 (right) shows the distribution of $\langle R^{e}_{0}\rangle$ versus $\langle R^{e}_{1}\rangle$ over the set of all videos. The rate of chats induced by videos is typically one-fourth the rate of spontaneous chats. In summary, this highlights that live chat discussions elicit emotional expressions more effectively through social dynamics compared to passive video content consumption. In Appendix E, we further support these findings with additional robustness checks, varying different aspects of our study design, modeling selected subsets of the data sample and emotion categories. Below, we further show that there is some heterogeneity across video categories, which, however, does not affect the main conclusions of our study.
@@
Across our 100 bootstrapped estimates, we estimate the spectral radius to be $0.80\pm 0.008$, which places the system well below criticality. This results in an expected cascade size of $5.0\pm 0.2$, meaning that each exogenous emotion, on average, generates a cascade of five total events before activity dies out. Thus, while the system exhibits substantial self-excitation, it remains far from the near-critical regime. A similar pattern holds at the level of individual emotions, where the rate of endogeneity for each emotion type $f$ is given by $\sum_{e}\alpha^{e,f}$.
@@
Variation Across Video Categories
Our main analysis examines general emotional and behavioral patterns among online users by compiling 92,412 YouTube live chat messages across 27 topics. The susceptibility of emotion dynamics to video influence and peer interactions is potentially subject to the type of video and the audience community. Here, we conduct sub-sample analyses on representative video types to identify heterogeneity in the evolution of online emotion dynamics. We focus on 4 video categories with keywords “politics”, “sports”, “live”, and “podcast”, with sample sizes 3,043, 11,172, 4,156, and 5,836 live chat messages, respectively. We report the model parameter estimates and relative ratios of different sources of emotion intensity in Appendix E.
Across different video categories, we observe several consistent patterns that align with our main findings: joy remains the most contagious emotion and exhibits the highest spontaneous intensity. Emotion dynamics are largely dominated by endogenous intensities except for the emotions of “surprise” and “fear”, which are shown to be more responsive to external stimuli, particularly for sport events and podcast shows. We also identify distinct patterns associated with specific video types and audience characteristics. The cross-triggering effects of anger and disgust are particularly potent for political videos, which contribute to moral outrage commonly observed in the political context. Unlike our main result (Figure 3), the content of political videos has a comparatively larger influence on the emotion dynamics of viewers, consistent with their provocative nature. In podcast and live-themed shows, surprise demonstrates a particularly long-lasting direct effect. Moreover, surprising content in podcast videos elicits joyful reactions. These findings validate our methodology by demonstrating its ability to disentangle the distinct channels through which video content and audience characteristics shape emotional dynamics.
In Appendix E, we show that our results remain robust when using alternative parameterizations of video influence, varying data preprocessing steps, excluding videos that primarily feature interactions between the audience and the video content, or modeling a reduced set of emotions.